# Electrocardiogram Instruction Tuning for Report Generation

### v2

@@ Abstract
Electrocardiogram (ECG) serves as the primary non-invasive diagnostic tool for cardiac conditions monitoring, are crucial in assisting clinicians. Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is not only time-consuming but also requires clinical expertise. To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions. To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets. Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open source LLMs, using more than 800,000 ECG reports. MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation. These findings emphasize the efficacy of our MEIT [^1] All data and code will be available upon acceptance. framework and its potential for real-world clinical application.
Machine Learning, ICML
[…]
${}^{\text{1}}$he Ohio State University ${}^{\text{2}}$niversity of North Carolina at Chapel Hill ${}^{\text{3}}$mperial College London
${}^{\text{4}}$ King’s College London  ${}^{\text{5}}$ ${}^{\text{6}}$
$*{}^{*}$orrespondence to: Zhongwei Wan<wan.512@osu.edu>Mi Zhang<mizhang.1@osu.edu>
${}^{\text{\textdagger}}$ore contribution  ${}^{\ddagger}$ore advising
@@ 1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports. Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, artificial intelligence (AI) models have been developed to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b). Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area. Unlike other AI-empowered medical report generation applications (e.g., radiology report), the primary challenges for ECG report generation stem from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. This difference in content and semantic interpretation between imaging and signal data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
@@
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions. Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space, enabling LLMs to understand signals for report generation without introducing additional training parameters in the attention layer.
In addition to the ECG report generation approach within MEIT, we introduce a comprehensive benchmark for ECG report evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation. Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs. The results demonstrate (1) the superior performance of MEIT in ECG report generation and the effective learning and alignment of ECG representations; (2) the effective transferability of LLMs under the MEIT framework in domain transfer scenarios.
@@
To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs. This framework incorporates a lightweight, attention-based fusion module effective across various LLM models. Furthermore, we establish a new benchmark for ECG report generation, which contains three aspects of the evaluation tasks. Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot test and robustness against data perturbations. The MEIT framework paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
@@ 2 Preliminaries
Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time. An ECG recording typically contains a 12-lead multivariate time series, which act as a 12-dimensional sequence of embeddings. The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function. ECG leads can be categorized into 6 limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and 6 precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views. We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the numbers of leads, $T$ is signal length. Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description. Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$. More visualization details can be found in the Appendix.
@@ 3 MEIT
Figure 1 illustrates the proposed MEIT framework. We begin by describing the curation process for ECG instruction data. We then delve into the architecture of the multimodal LLM designed for ECG report generation. Lastly, we describe the approach used to generate ECG reports with instruction tuning.
@@ 3.1 ECG Instruction Data Curation
Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt. For instance, the prompt might be, “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ”. To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$. In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts. To address this challenge, we first manually design some prompt samples, then utilize GPT-4 (Achiam et al., 2023) to generate a set of prompts by rephrasing, as shown in Figure 1. Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set, and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response. This approach ensures that the generated response not only conveys the same meaning as the ground truth but also remains adaptable to different instruction prompts. Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG (Gow et al., ) dataset that contains 80K annotated data, and a 20K dataset PTB-XL (Wagner et al., 2020).
@@ 3.2 ECG Report Generation Model
In MEIT framework, the multimodal ECG report generation model aims to decode ECG signals end-to-end and accurately generate ECG reports. The architecture is illustrated in Figure 2. Specifically, the report generation model aims to encode the ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate them with the language embeddings with modality alignment, and then autoregressively generate the ECG report. In the following, we detail each component of the report generation model.
ECG Encoder. Since ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone. Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of some temporal convolution blocks to encode each ECG signal into embeddings. Specifically, each temporal convolution block is composed of several 1-D convolution layers, batch normalization layers, and ReLU activation layers followed by average pooling. To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear up-to-down projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
@@
where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs. Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period.
ECG Modality Alignment. To guide the LLMs in aligning ECG signal data with corresponding textual outputs, we introduce an ECG modality alignment strategy. This approach is detailed in Figure 2. Specifically, given the ECG embeddings $\mathbf{Z}_{e}$, the alignment strategy aims to incorporate $\mathbf{Z}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task. Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022) ,and G-MAP (Wan et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
@@
here, $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone. We replicate $\mathbf{H}_{e}$ for each head $k$ times, merging the ECG and language features in the sequence dimension. This is achieved through a shared projection of keys and values for each pattern. The fusion is then articulated as:
@@
where $\mathbf{Q}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{q,j}$, $\mathbf{K}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{k,j}$, and $\mathbf{K}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{k,j}$, with a similar notation for $\mathbf{V}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{v,j}$ and $\mathbf{V}_{t,j}=\mathbf{H}_{t}^{i}\mathbf{W}_{v,j}$. Concatenation is denoted by []delimited-[][$\cdot$][ ⋅ ], and $\mathbf{K}_{m,j}$ and $\mathbf{V}_{m,j}$ symbolize the amalgamated features of query and key. $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively. Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality.
@@ 3.3 ECG Instruction Tuning
As described in Section 3.1, we have converted ECG-text pairs into a chat-bot style instruction format: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$}</s>. During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$. To save computational resources and accelerate the convergence of instruction tuning, we use the LoRA (Hu et al., 2021) adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone. Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:

### v3

# MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation

@@ Abstract
Electrocardiogram (ECG) is the primary non-invasive diagnostic tool for monitoring cardiac conditions and is crucial in assisting clinicians. Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is time-consuming and requires clinical expertise. To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions. To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets. Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open-source LLMs using more than 800,000 ECG reports. MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation. These findings emphasize the efficacy of our MEIT framework and its potential for real-world clinical application.
@@ 1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which include electrocardiogram (ECG) recordings and clinical reports. Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, AI models have been developed to facilitate ECG data analysis for the task of classification [1, 2, 3, 4]. Despite these efforts, the automatic generation of reports from ECG recordings still needs to be explored. Unlike other AI-empowered medical report generation applications (e.g., radiology reports), the primary challenge for ECG report generation stems from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. The difference in the content and semantic interpretation between imaging and ECG data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
@@
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG recordings and human instructions. Inspired by the versatility of LLMs [5, 6, 7, 8, 9] in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG recordings, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space. This enables LLMs to understand ECG signals for report generation without introducing additional training parameters in the attention layer. In addition to the ECG report generation approach, we introduce a comprehensive benchmark for ECG report generation evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation. Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs. The results demonstrate (1) the superior performance of MEIT in ECG report generation and the effective learning and alignment of ECG representations; (2) the effective transferability of LLMs under the MEIT framework in domain transfer scenarios.
@@
To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs. This framework incorporates a lightweight, attention-based fusion module across various LLM models. Furthermore, we design a new benchmark for ECG report generation, which contains three evaluation tasks. Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot tests and robustness against data perturbations. MEIT paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
@@ 2 Preliminaries
Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time. An ECG recording typically contains a 12-lead multivariate time series, which acts as a 12-dimensional sequence of embeddings. The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function. ECG leads can be categorized into six limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views. We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the number of leads, and $T$ is signal length. Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description. Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$. More details on visualization can be found in the appendix A.5.
@@ 3 MEIT
Figure 2 illustrates the proposed MEIT framework. We begin by describing the curation process for ECG instruction data. We then delve into the architecture of the multimodal LLM designed for ECG report generation. Lastly, we describe the approach to generating ECG reports with instruction tuning.
@@ 3.1 ECG Instruction Data Curation
Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt. For instance, the prompt can be “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ". To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$. In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts. To address this challenge, we manually design some prompt samples, then utilize GPT-4 [5] to generate a set of prompts by rephrasing, as shown in Figure 2. Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response. This approach ensures that the generated response conveys the same meaning as the ground truth and remains adaptable to different instruction prompts. Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG [10] dataset that contains 80K annotated data and a 20K dataset PTB-XL [11]. The ECG instruction data samples are shown in Appendix A.5.
@@ 3.2 ECG Report Generation Model
In MEIT, the multimodal ECG report generation model decodes ECG signals end-to-end to generate ECG reports. The architecture is illustrated in Figure 2. Specifically, the report generation model directly encodes an entire ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate it with the language embeddings with modality alignment, and then autoregressively generate the ECG report. Next, we detail each component of the report generation model.
ECG Encoder. Since the ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone. Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode each ECG signal into embeddings. Specifically, each temporal convolution block comprises several 1-D convolution layers, batch normalization layers, and ReLU activation layers, followed by average pooling. To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
@@
where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs. Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period. More details about ECG Encoder are illustrated in Appendix A.3.
ECG Modality Alignment. We introduce an ECG modality alignment strategy to guide the LLMs in aligning ECG signal data with corresponding textual outputs. This approach is detailed in Figure 2. Specifically, given the ECG embeddings $\mathbf{H}_{e}$, the alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task. Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo [12], Memorizing Transformer [13] , and G-MAP [14], or Q-former in BLIP-2 [15] that require the additional trainable parameters and design complex modules for two-stage training, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
@@
where $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone. We replicate $\mathbf{H}_{e}$ for each head $k$ times, merging the ECG and language features in the sequence dimension. This is achieved through a shared projection of keys and values for each pattern. The fusion is then articulated as:
@@
where $\mathbf{Q}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{q,j}$, $\mathbf{K}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{k,j}$, and $\mathbf{K}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{k,j}$, with a similar notation for $\mathbf{V}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{v,j}$ and $\mathbf{V}_{t,j}=\mathbf{H}_{t}^{i}\mathbf{W}_{v,j}$. Concatenation is denoted by []delimited-[][$\cdot$][ ⋅ ], and $\mathbf{K}_{m,j}$ and $\mathbf{V}_{m,j}$ symbolize the amalgamated features of query and key. $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively. Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality. More comparisons about ECG modality alignment and other fusion approaches are illustrated in Appendix LABEL:appendix:_ECG_align_ablation.
@@ 3.3 ECG Instruction Tuning
As described in Section 3.1, we have converted ECG-text pairs into a chat-bot style instruction format: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$}</s>. During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$. To save computational resources and accelerate the convergence of instruction tuning, we use LoRA [16] adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone. Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:
@@ 4 Benchmarking ECG Report Generation
In this section, we create a benchmark to evaluate the ECG report generation approach. Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation, (2) the zero-shot ability under domain transfer, (3) the robustness of the model against perturbed signals.
4.1 Datasets
PTB-XL. The PTB-XL dataset [11] comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients. Each ECG is accompanied by a corresponding report. The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds. We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively. The human experts double-check all samples in the test set to ensure data quality. As mentioned in Sec 3.1, we reformulate the dataset into the instruction data format.
MIMIC-IV-ECG. The MIMIC-IV-ECG dataset [10] is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects. Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds. The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio. Likewise, we reconstruct this dataset into an ECG instruction data template.
4.2 Evaluation Metrics
We evaluate models using nine metrics: BLEU 1-4 [17], METEOR [18], ROUGE 1-2 and L [19], CIDEr-D [20], and BERTScore [21]. BLEU and METEOR assess machine translation quality, focusing on accuracy and fluency. ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps. CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
4.3 Tasks
Quality Report Generation. This task aims to assess report quality after ECG instruction tuning using 10% of MIMIC-IV-ECG and PTB-XL datasets as the test set. The evaluation examines how closely generated reports match the original’s structure and meaning, considering various instructions and ECG inputs. We analyze metrics like BLEU-1 to 4, METEOR, ROUGE 1, 2, L, CIDEr-D, and BERTScore.
Zero-shot Generalizability. To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG. Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set. It’s important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents—Europe and the United States, respectively—utilizing varied devices and from distinct hospitals, across different time periods. Therefore, we consider these datasets to represent two separate domains. This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively. We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
5 Experiments and Evaluation
5.1 Experimental Settings
Models. We use nine LLMs based on the peft [^1] https://github.com/huggingface/peft library, which directly supports LoRA [16] to construct the multimodal ECG report generation model described in Section 3.2. These models include GPT-Neo [22], GPT-NeoX [23], GPT-J [24], BLOOM [25], OPT [26], LLaMA-1 [27], LLaMA-2 [6], Mistral [28], and Mistral-Instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with